1) When Jasper LLC produces 500 units, the total cost is $8750; what is the average total cost?

  1. 21.20
  2. 21.00
  3. 17.50
  4. 19.21
  5. none of the above

2) Jasper LLC's total cost is $10500 and the average total cost is $17.50; how many units are they producing?__________

3) When Jasper LLC produces Q=700, ATC=17.50; what is TC? ______

4)Refer to questions #5, #6, and #7. Jasper LLC is experiencing

  1. economies of scale.
  2. constant returns to scale.
  3. diseconomies of scale
  4. a labor shortage

5) Suppose a consumer of oranges and tangerines is currently in equilibrium. If the price of tangerines falls, the consumer will attempt to restore equilibrium by:

A) buying more oranges

B )buying fewer tangerines

C) buying more tangerines

D) A and B

E) None of the above
Answer #1

1) ATC = TC/Q

= 8750/500 = 17.5

2) TC=10500

ATC=17.5

Q = 10500/17.5 = 600

3) Q=700

ATC=17.5

TC = ATC*Q = 700*17.5 = 12250

4) Since ATC remains the same, Jasper LLC is experiencing:-

Constant returns to scale.